another Suppose this player gets to shoot three free throws Find the probability that he misses all three consecuive free twows Round to the nearest en-thousandth OA. 0.0005 OB. 09995 O C. 0.7787 O D. 0.2213
Answer #1

Required probability = Probability that Player misses consecutively all three free throws

We are given,

Probability that player makes a single free throw = 92% = 0.92

Complement of the above probability will be :

Probability that player does not make(misses) a single free throw = 1 - 0.92 = 0.08

Probability that Player misses consecutively all three free throws =

= ( Probability that player does not make(misses) a single free throw )3

= ( 0.08)3 = 0.0005

Correct option : A. 0.0005
Answer #2

subtract 


i did

 100-91=8

multiply 

         3

(0.08)


round 

0.000729


correct answer

0.0007

answered by: annajacob123